OpenStudy (anonymous):

the height of a cone is two times its base diameter what is the volume of the cone in terms of its base radius.

OpenStudy (anonymous):

please help me

OpenStudy (cwrw238):

the formula for the volume of a cone is (1/3)pi r^2 h where r = radius of the base and h = height of the cone if h = 2*d and d = 2r then h = ? can you work that out

OpenStudy (cwrw238):

when yiu have found h in terms of r just substitute this into the formula for the volume

OpenStudy (anonymous):

thank you let me try